### v2

# Possible Unconventional Surface Superconductivity in the Half-Heusler YPtBi

Abstract
We report an extensive study of the noncentrosymmetric half-Heusler superconductor YPtBi, revealing an unusual relation between bulk superconductivity and the possible appearance of surface superconductivity on the (111) oriented surface, at temperatures up to 3 times the bulk transition temperature. Transport measurements confirmed the low carrier density of the material and its bulk superconducting transition, which was also observed in ac susceptibility through mutual inductance (MI) measurements. However, a weak signature of superconductivity in the MI measurements appeared much above the bulk transition temperature, which was further observed in scanning tunneling spectroscopy, pointing to a possible surface superconducting state. Polar Kerr effect measurements suggest that while the bulk superconductor may exhibit an unusual nodal superconducting state, only the surface state breaks time reversal symmetry. Complementary tunneling measurements on LuPtBi are used to establish the observations on YPtBi, while density-functional theory calculations may shed light on the origin of this unusual surface state.
Topological superconductivity can be realized by inducing superconductivity at the surface of a three dimensional topological insulator [1]. This can be achieved through interface engineering, or by using the surface states of a bulk superconductor [4, 3, 2]. These mechanisms generally lead to time-reversal invariant topological states [1], and their critical temperature is determined by that of the bulk superconductor. An alternative approach, which expands the limits of these techniques, is utilizing flat surface bands which can support superconductivity without a proximity effect. Such superconducting surface states have been observed in KZnBi [5] and t-PtBi2 [6]. However, the properties of these states, particularly the origin and nature of their superconductivity, are not fully understood.
Here, we explore the possible formation of a superconducting surface state at the (111) surface of the half Heusler compound YPtBi through extensive measurements that include transport, ac susceptibility measurements through mutual inductance, scanning tunneling microscopy and spectroscopy and polar Kerr effect. While robust bulk superconductivity appears below , we find that superconductivity at the (111) surface persists to temperatures . Furthermore, our Kerr effect measurements suggest that unlike the bulk, this superconducting surface state breaks time reversal symmetry (TRS), thus revealing a delicate interplay between the surface and bulk states. We complement these measurements with scanning tunneling spectroscopy and mutual inductance of (111) oriented LuPtBi to substantiate our arguments. To further support our experimental findings, we performed ab-initio calculations on the Bi, Lu and Pt terminations of the (111) direction, finding that a Bi-termination surface exhibits a van-Hove singularity (VHS) near the Fermi energy, which could be a source for for the superconducting surface state. A recent theoretical study [7] explored a similar surface band structure, finding a possible domain for the onset of chiral topological surface superconductivity ahead of the bulk transition, in agreement with our results. Our results suggest a new platform for 2D chiral superconductors in which a surface state condenses into a TRS-breaking superconducting state independently from the bulk.
In a half-Heusler compound, three metallic elements are arranged in interpenetrating fcc lattices ( structure) [8]. The choice of elements, with thousands of possible combinations, enables control over various properties in the resulting compounds: structures can be metallic, semiconducting, magnetic or non-magnetic, superconducting, and they can have topological surface states [9, 10]. Of these materials, YPtBi and LuPtBi have the largest inverted gaps [11], leading to topologically protected surface states [12]. Furthermore, YPtBi and LuPtBi are bulk superconductors, with critical temperatures of 0.8 K and 1 K respectively [13, 14]. The bulk superconductivity should be a combination of odd and even parity states, due to the lack of inversion symmetry [15]. While much work was done to explore the bulk superconductivity in these materials [13, 14, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], studies of the surface states and their propensity to unconventional superconductivity have been limited [7].
We studied single crystals of LuPtBi and YPtBi that were grown by self-flux solution growth method, with Bi used as the flux medium [31, 32, 33]. The crystal structure and composition were verified using dispersive X-ray analysis, Laue X-ray diffraction and X-ray Photoelectron Spectroscopy [33]. Transport measurements confirmed the low carrier densities, /cm3, and relatively high mobilities, typically cm2/V-s at helium temperature. Sharp bulk superconductivity was verified with both resistive transition and two-coils mutual inductance (MI) measurements. The scanning Tunneling Microscopy (STM) and spectroscopy (STS) measurements were performed after removing the top few surface layers with the STM tip itself. The details of this surface cleaning procedure are described in the Supplemental Material [33], and are based on a similar procedure previously used to expose superconducting regions in vacuum-cleaved YPtBi crystals [34].
Fig. 1 shows representative base temperature (350 mK) spectra of pristine surface and tip-cleaned surface for YPtBi and LuPtBi samples. While the tip-cleaned surface improved the spectra, particularly in reducing the observed zero bias conductance to very low values, the pristine surface spectra pointed to a similar large gap. The observed spectra are rounded, even well below temperatures corresponding to the observed large gap. Such a behavior could be because of a nodal gap structure or high effective temperature due to RF heating. The observation of a flat gap (Fig. 1b) rules these options out. Another possibility is that a thin metallic layer between the superconducting state and the tip depresses the gap by proximity effect (Fig. 1c) [37]. In this scenario, the variability in the flatness of the gap and the density of states at zero bias [33] is explained by the inhomogeneous metallic film that is created by the tip cleaning. Figure 1c illustrates the proposed effect of the current pulse. Before the pulse, a thin disordered metallic layer covers the surface, and the pulse removes some, but not all, of this layer, resulting in a variable metallic barrier, which rounds or supresses the gap at different spots.
The large gaps observed at base temperature indicate that there is a superconducting system in both LuPtBi and YPtBi which involves a transition temperature higher than the bulk. Indeed, temperature dependence of tunneling spectra of both materials indicates gaps that persist much above the bulk . Fig. 2a shows a typical set of spectra taken on YPtBi at different temperatures. To estimate the gap, we fitted the spectra to the Dynes formula [38, 33]. The temperature was left as a fitting parameter, with a higher effective temperature required to account for the broadening of the spectra. Figure 2b shows the extracted gaps plotted against the measured temperature. The gap at our base temperature was 0.4 meV. Assuming yields . We note that while the spectra taken following the tip-cleaning procedure can be fitted to the Dynes formula, this procedure fails for the spectra taken on the pristine surface, due to the large residual density of states. This is consistent with the presence of a thin metallic barrier between the tip and the superconducting layer, as has been observed and modeled for thin metal-supercodnuctor interfaces [37]. This metallic barrier is then removed by the pulsing procedure.
The tip cleaning procedure could change the composition on the surface or the lattice structure. Indeed, the procedure used in Ref. [34] generated craters around the pulse location, in which a polycrystalline system was formed. The confinement of superconductivity to these craters could result in an increaed gap. However, the gaps we observed before and after the tip-cleaning procedure were of similar sizes, suggesting that the cleaning does not induce the larger gaps. Furthermore, the topography of our samples following the pulsing procedure did not show craters, and spectra did not systematically vary with the distance from the pulse [33].
A tip-induced enhancement of superconductivity is further ruled out by measurements of the ac susceptibility of both materials, using a mutual inductance technique [40]. These crystals were not treated before the measurements. Both materials showed a large increase of the diamagnetic response at their respective bulk transition temperatures (Fig. 3), in agreement with previous reports [13, 14]. However, a careful measurement of YPtBi revealed a small increase, 0.5% of the change due to the bulk transition, which onsets at 3 K (Fig. 3a). LuPtBi showed a similar change, at around 5 K, over a background that is caused by thermal expansion (Fig. 3b and [33]). This may indicate that a small volume fraction of both samples turns superconducting at a higher temperature than the bulk.
The observation of a similar transition temperature for the vanishing of the superconducting gap in tunneling, and for the onset of an inductive signal in mutual inductance points to an intrinsic origin of superconductivity, that onsets at a temperature much above the bulk . We envision two possible origins for the observations. First, given our inability to make clean, atomically flat cleaves that are known to expose the (111) plane, the signals may be attributed to another compound, created during the surface cleaning, a combination of the HH elements, possibly with contaminants, particularly oxygen. Superconducting compounds that may be formed this way are listed in Table 1. While the of both amorphous and polycrystalline Bi are suitable, such compounds on the surface would generate similar transitions for both LuPtBi and YPtBi, inconsistent with the difference in gap sizes and onset temperatures. We further note that LuPtBi and YPtBi are very similar chemically and structurally (lattice constants Å and Å, respectively [11]), suggesting that a strained inclusion of a different compound on the surfaces would yield a similar gap. Moreover, the similarity in the behavior of LuBi and YBi may rule this possibility out as well, particularly since they require pressure to exhibit superconductivity.
Compound | Bi | a-Bi | p-Bi | BiPt | Bi2Pt | YBi | LuBi |
P (GPa) | amb. | amb. | amb. | amb. | 5-6 | ||
5-62 | 6.13 | 1.34 | 25 | 56 | 4-7.57 |
The alternative explanation for the observed higher in the two compounds is the creation of a chiral surface state. Particularly, since both LuPtBi and YPtBi are known to be topological with a large band inversion [11], it was suggested that superconductivity associated with this material is topological and may break TRS [22, 23]. To test this possibility, we performed high resolution polar Kerr effect measurements on YPtBi using a zero-area-loop Sagnac interferometer (ZALSI) [35, 36] (see supplementary material [33]).
We cooled down the sample to 300 mK from a high temperature in the presence of a small external magnetic field, and then removed the field and measured the Kerr signal while warming up the sample (Fig. 4a). For small training fields ( Oe) the initial signal (at 300 mK) following the removal of the magnetic field was very small, consistent with zero (within our resolution of rad). However, increasing the temperature from base temperature, a finite Kerr signal of order rad appeared above the bulk . The signal was independent of the field strength and changed sign upon changing the direction of the training field (Fig. 4b and [33]). These features indicate a TRS breaking state, and are generally similar to those observed in chiral superconductors [49, 50]. Note that a TRSB state further restricts the association of the signals observed using mutual inductance and STM with a different chemical structure of the surface. In particular, amorphous and polycrystalline Bi are not chiral superconductors, and therefore will not give rise to a finite Kerr effect. The contribution from vortices [33] trapped by the field-cooling procedure is too small to account for the observed signals.
With strong structural and electronic structure similarities to GdPtBi, we estimate the optical penetration depth of YPtBi at nm [51]. Thus, the Kerr signal can have both bulk and surface contributions. In particular above the bulk the observed Kerr signal may be a result of TRS breaking associated with the surface superconductivity - a favorable possibility since Kerr effect persists very close to the surface , or it may be associated with vestigial order [52, 53, 54, 55] originating from the bulk superconducting state. However, the suppression of the Kerr rotation below , as well as the nodes in the gap and analysis of the possible spin and orbital pairing states [23] suggest that the bulk does not break TRS.
While low field training (i.e. training in Oe) yields similar onset and size of Kerr signal, which follows the direction of the training field, cooling the sample at higher magnetic field to base temperature (0.3 K), turning the field to zero and measuring while warming up the sample in zero magnetic field yields a signal that is zero within our resolution of rad (Fig. 5a). However, cooling the sample at higher magnetic field to base temperature (0.3 K), but on the way turning the magnetic field to zero at K, then measuring the Kerr effect at zero field from base temperature up yields results similar to the low field training experiments (Fig. 5b,c).
Such a peculiar effect may point to a delicate interplay between the vortex state that appears in the bulk material upon cooling in a field much below the bulk , and the TRSB surface state. For example, cooling at lower field and removing the field at low temperatures may result in low vortex occupancy due to weak surface and geometrical barriers (note that Oe, while the thermodynamic critical field is Oe for the bulk superconducting state of these materials). Above a field of order Oe, enough vortices may remain in the bulk, which upon crossing the bulk Tc will move out of the sample, thus possibly “scrambling” the single domain nature of the TRS-breaking state of the surface superconductivity. Thus, the magnetization of the surface state interacts with a conventional bulk superconducting order parameter. This is similar to magnetic superconductors, where the formation and motion of vortices can change the ferromagnetic order parameter [56].
Taken together, the STM, mutual inductance and PKE results indicate that the (111) surfaces of YPtBi and LuPtBi host a superconudcting state which breaks TRS. To understand the origin of this state, we studied the surface band structure of LuPtBi (see Supplementary Information [33]). In Bi-terminated (111) surfaces, we found a van-Hove signularity close to the -point, resulting in an enhanced density of states near the Fermi level, which could give rise to the enhanced surface . The calculated band structure of the Lu terminated surface did not show van-Hove singularities. Schwemmer et al. [7] have recently reported a detailed study of the surface band structure and possible superconductivity of the Bi-terminated (111) surface of LuPtBi. While the gross features of their surface band structure are similar to our calculation, they also revealed flat-band regions near the Fermi energy, and calculated the propensity to a superconducting instability. In particular, they argue that the transition temperature of the surface state is independent of the bulk transition, and exhibit vestigial time-reversal symmetry breaking in it fluctuation regime. Since inversion symmetry is broken, simple singlet and/or triplet states were dismissed and pairing was tested for pairs of electrons that are connected by the TR operator. The result has been a pair-amplitude that transforms under the E irreducible representation of the surface point group (relevant to the (111) crystal orientation) in the presence of strong Kane-Mele type and Rashba type spin-orbit coupling [57]. Furthermore, they found that a chiral superconducting state, which spontaneously breaks TRS, is energentically favorable. While we focused on the (111) oriented surface, surface bands close to the Fermi energy that originate from dangling bonds are generic in the half-Heuslers along other surfaces, including the (001) [12, 58] and (120) [30]. It is therefore possible that the surface superconductivity we report here could also exist on other surfaces, depending on the filling and flatness of the surface bands.
In conclusion, we presented a comprehensive study of the (111) surface of the half-Heusler alloys YPtBi and LuPtBi. While a bulk of 0.8 K is observed for both systems, we find a much larger corresponding gap in our tunneling studies with ’s of 2.3 K and 6.5 K for YPtBi and LuPtBi, respectively. Mutual inductance measurements of YPtBi further revealed a small diamagnetic response that onsets around . Polar Kerr effect on YPtBi indicates that TRS is broken near or at . Together, our data suggest that chiral surface superconductivity precedes bulk condensation in these materials, a hypothesis which is strongly supported by recent calculations [7]. While additional work is required to determine the order parameter of these states, our results open a new direction to engineering topological superconducting states in half Heusler compounds. Although our study was conducted on single crystals, Kim et al. [59] recently demonstrated epitaxially growth of YPtBi films, which can also be encapsulated by an alumina film, preventing degradation of the surface. This system is therefore highly promising for quantum device applications.
Acknowledgements.
We thank Ronny Thomale and Jörg Schmalian for illuminating discussions. This work was supported by the U. S. Department of Energy (DOE) Office of Basic Energy Science, Division of Materials Science and Engineering at Stanford under contract No. DE-AC02-76SF00515 and by the National Science Foundation Award Number: 2307132. EP was partially supported by the Koret Foundation.References
- [1] L. Fu, and C. L. Kane, “Superconducting Proximity Effect and Majorana Fermions at the Surface of a Topological Insulator,” Phys. Rev. Lett. 100, 096407 (2008)
- [2] A.K. Nayak, A. Steinbok, Y. Roet, J. Koo, G. Margalit, I. Feldman, A. Almoalem, A. Kanigel, G. Fiete, B. Yan, Y. Oreg, N. Avraham and H. Beidenopf, “Evidence of topological boundary modes with topological nodal-point superconductivity,” Nat. Phys. 17, 1413–1419 (2021).
- [3] O.J. Clark, M.J. Neat, K. Okawa, L. Bawden, I. Marković, F. Mazzola, J. Feng, V. Sunko, J.M. Riley, W. Meevasana, J. Fujii, I. Vobornik, T.K. Kim, M. Hoesch, T. Sasagawa, P. Wahl, M.S. Bahramy, and P.D.C. King,“Fermiology and Superconductivity of Topological Surface States in PdTe2,” Phys. Rev. Lett. 120, 156401 (2018).
- [4] P. Zhang, K. Yaji, T. Hashimoto, Y. Ota, T. Kondo, K. Okazani, Z. Wang, J. Wen, G.D. GU, H. Ding, and S. Shin, “Observation of topological superconductivity on the surface of an iron-based superconductor,” Science 360, 182-186 (2018).
- [5] J. Song, S. Kim, Y. Kim, H. Fu, J. Koo, Z. Wang, G. Lee, J. Lee, S.H. Oh, J. Bang, T. Matsushita, N. Wada, H. Ikegami, J.D. Denlinger, Y.H. Lee, B. Yan, Y. Kim, and S.W. Kim, “Coexistence of Surface Superconducting and Three-Dimensional Topological Dirac States in Semimetal KZnBi,” Phys. Rev. X 11, 021065 (2021).
- [6] Sebastian Schimmel, Yanina Fasano, Sven Hoffmann, Joaquín Puig, Grigory Shipunov, Danny Baumann, Saicharan Aswartham, Bernd Büchner, and Christian Hess, “High- surface superconductivity in topological Weyl semimetal t-PtBi2,” arXiv:2302.08968 [cond-mat.supr-con] (2023).
- [7] Tilman Schwemmer, Domenico Di Sante, Jörg Schmalian, Ronny Thomale, “Chiral surface superconductivity in half-Heusler semimetals,” arXiv:2212.09786 [cond-mat.supr-con] (2022).
- [8] Tanja Graf, Claudia Felser, and Stuart S.P. Parkin, “Simple rules for the understanding of Heusler compounds,” Prog. Solid. State Ch. 39, 1-50 (2011).
- [9] Kaustuv Manna, Yan Sun, Lukas Muechler, Jürgen Kübler, and Claudia Felser, “Heusler, Weyl and Berry,” Nat Rev Mater 3, 244–256 (2018).
- [10] Sheron Tavares, Kesong Yang, and Marc A. Meyers, “Heusler alloys: Past, properties, new alloys, and prospects,” Prog. Mater. Sci. 132, 101017 (2023).
- [11] S. Chadov, X. L. Qi, J. Kübler, G. H. Fecher, C. Felser, and S. C. Zhang, “Tunable multifunctional topological insulators in ternary Heusler compounds,” Nat. Mater. 9, 541 (2010).
- [12] Z. K. Liu, L. X. Yang, S.-C. Wu, C. Shekhar, J. Jiang, H. F. Yang, Y. Zhang, S.-K. Mo, Z. Hussain, B. Yan, C. Felser, and Y. L. Chen, “Observation of unusual topological surface states in half-Heusler compounds LnPtBi (Ln=Lu, Y),” Nat. Commun 7, 12924 (2016).
- [13] N. P. Butch, P. Syers, K. Kirshenbaum, A. P. Hope, and J. Paglione, “Superconductivity in the topological semimetal YPtBi,” Phys. Rev. B 84, 220504 (2011).
- [14] F. F. Tafti, Takenori Fujii, A. Juneau-Fecteau, S. René de Cotret, N. Doiron-Leyraud, Atsushi Asamitsu, and Louis Taillefer, “Superconductivity in the noncentrosymmetric half-Heusler compound LuPtBi: A candidate for topological superconductivity,” Phys. Rev. B 87, 184504 (2013).
- [15] M Smidman, M. B. Salamon, H. Q. Yuan, and D. F. Agterberg, “Superconductivity and spin–orbit coupling in non-centrosymmetric materials: a review,” Rep. Prog. Phys. 80, 036501 (2017).
- [16] T. V. Bay, T. Naka, Y. K. Huang, and A. de Visser, “Superconductivity in noncentrosymmetric YPtBi under pressure,” Phys. Rev. B 86, 064515 (2012).
- [17] T.V. Bay, M. Jackson, C. Paulsen, C. Baines, A. Amato, T. Orvis, M.C. Aronson, Y.K. Huang, A. de Visser, “Low field magnetic response of the non-centrosymmetric superconductor YPtBi,” Solid State Comm. 183, 13 (2014).
- [18] Markus Meinert, “Unconventional Superconductivity in YPtBi and Related Topological Semimetals,” Phys. Rev. Lett. 116, 137001 (2016).
- [19] Orest Pavlosiuk, Dariusz Kaczorowski, and Piotr Wiśniewski, “Superconductivity and Shubnikov–de Haas oscillations in the noncentrosymmetric half-Heusler compound YPtBi,” Phys. Rev. B 94, 035130 (2016).
- [20] C. Timm, A. P. Schnyder, D. F. Agterberg, and P. M. R. Brydon, “Inflated nodes and surface states in superconducting half-Heusler compounds,” Phys. Rev. B 96, 094526 (2017).
- [21] Hao Yang, Jiabin Yu, Stuart S.P. Parkin, Claudia Felser, Chao Xing Liu, and Binghai Yan, “Prediction of Triple Point Fermions in Simple Half-Heusler Topological Insulators,” Phys. Rev. Lett. 119, 136401 (2017).
- [22] Lucile Savary, Jonathan Ruhman, Jörn W. F. Venderbos, Liang Fu, and Patrick A. Lee, “Superconductivity in three-dimensional spin-orbit coupled semimetals,” Phys. Rev. B 96, 214514 (2017).
- [23] Hyunsoo Kim, Kefeng Wang, Yasuyuki Nakajima, Rongwei Hu, Steven Ziemak, Paul Syers, Limin Wang, Halyna Hodovanets, Jonathan D. Denlinger, Philip M. R. Brydon, Daniel F. Agterberg, Makariy A. Tanatar, Ruslan Prozorov, Johnpierre Paglione, “Beyond triplet: Unconventional superconductivity in a spin-3/2 topological semimetal,” Sci. Adv. 4, eaao4513 (2018).
- [24] Jörn W.F. Venderbos, Lucile Savary, Jonathan Ruhman, Patrick A. Lee, and Liang Fu, “Pairing States of Spin-3/2 Fermions: Symmetry-Enforced Topological Gap Functions,” Phys. Rev. X 8, 011029 (2018).
- [25] Qing-Ze Wang, Jiabin Yu, and Chao-Xing Liu, “Unconventional superconductivity and surface pairing symmetry in half-Heusler compounds,” Phys. Rev. B 97, 224507 (2018).
- [26] Bitan Roy, Sayed Ali Akbar Ghorashi, Matthew S. Foster, and Andriy H. Nevidomskyy, “Topological superconductivity of spin-3/2 carriers in a three-dimensional doped Luttinger semimetal,” Phys. Rev. B 99, 054505 (2019).
- [27] Hyunsoo Kim, Makariy A. Tanatar, Halyna Hodovanets, Kefeng Wang, Johnpierre Paglione, and Ruslan Prozorov, “Campbell penetration depth in low carrier density superconductor YPtBi,” Phys. Rev. B 104, 014510 (2021).
- [28] Hyunsoo Kim, Junhyun Lee, Halyna Hodovanets, Kefeng Wang, Jay D. Sau, and Johnpierre Paglione, “Quantum oscillations of the Fermi surface in the topological semimetal YPtBi,” Phys. Rev. Research 4, 033169 (2022).
- [29] Y.Z. Zhou, J. Chen, Z.X. Li, J. Luo, J. Yang, Y.F. Guo, W.H. Wang, R. Zhou, and Guo-ging Zheng, “Antiferromagnetic Spin Fluctuations and Unconventional Superconductivity in Topological Superconductor Candidate YPtBi Revealed by 195Pt-NMR,” Phys. Rev. Lett. 130, 266002 (2023).
- [30] J. C. Souza, M. V. Ale Crivillero, H. Dawczak-Debicki, Andrzej Ptok, P. G. Pagliuso, and S. Wirth, “Tuning the topological character of half-Heusler systems: A comparative study on YTBi (T=Pd,Pt),” Phys. Rev. B 108, 165154 (2023).
- [31] P. C. Canfield, J. D. Thompson, W. P. Beyermann, A. Lacerda, M. F. Hundley, E. Peterson, Z. Fisk, and H. R. Ott, “Magnetism and heavy fermion‐like behavior in the RBiPt series,” J. Appl. Phys. 70, 5800 (1991).
- [32] P. C. Canfield and Z. Fisk, “Growth of single crystals from metallic fluxes,” Philos. Mag. B 65, 1117 (1992).
- [33] Supplemental Information - URL to be determined by publisher.
- [34] Hongwoo Baek, Jeonghoon Ha, Duming Zhang, Bharath Natarajan, Jonathan P. Winterstein, Renu Sharma, Rongwei Hu, Kefeng Wang, Steven Ziemak, Johnpierre Paglione, Young Kuk, Nikolai B. Zhitenev, and Joseph A. Stroscio, “Creating nanostructured superconductors on demand by local current annealing,” Phys. Rev. B 92, 094510 (2015).
- [35] Jing Xia, Peter T. Beyersdorf, Martin M. Fejer, Aharon Kapitulnik, “Modified Sagnac interferometer for high-sensitivity magneto-optic measurements at cryogenic temperatures,” Applied Physics Letters 89, 062508 (2006).
- [36] Aharon Kapitulnik, Jing Xia, Elizabeth Schemm and Alexander Palevski, “Polar Kerr effect as probe for time-reversal symmetry breaking in unconventional superconductors,” New J. Phys. 11, 055060 (2009).
- [37] N. Moussy, H. Courtois, and B. Pannetier, “Local spectroscopy of a proximity superconductor at very low temperature,” Europhys. Lett., 55, 861 (2001).
- [38] R. C. Dynes, V. Narayanamurti, and J. P. Garno, “Direct Measurement of Quasiparticle-Lifetime Broadening in a Strong-Coupled Superconductor,” Phys. Rev. Lett. 41, 1509 (1978).
- [39] Toshiyuki Kita, Daisuke Akasako, Hironaru Murakami and Ryozo Aoki, “Gap Value Determinations on the Tunneling Spectrum,” in Advances in Superconductivity VII, K. Yamafuji et al. (eds.), (Springer-Verlag Tokyo 1995.)
- [40] Xinyang Zhang, Mark Zic, Dong Chen, Chandra Shekhar, Claudia Felser, Ian R. Fisher, and Aharon Kapitulnik, “Vortex phase diagram of kagome superconductor CsV3Sb5,” arXiv:2306.13297 [cond-mat.supr-con] (2023).
- [41] O. Prakash, A. Kumar, A. Thamizhavel, S. Ramakrishnan, “Evidence for bulk superconductivity in pure bismuth single crystals at ambient pressure,” Science 355, 52 (2017).
- [42] J. T. Chen, T. T. Chen, J. D. Leslie and H. J. T. Smith, “Strong-Coupling Superconductivity in Amorphous Bismuth,” Physics Letters A 25, 679 (1967).
- [43] V. I. Petrosyan, V. N. Molin, O. I. Vasin, P. A. Skripkina, S. I. Stenin, and E. G. Batyev, “Superconductivity in polycrystalline bismuth films,” Zh. Eksp. Teor. Fiz. 66, 995-100 (1974).
- [44] Mingliang Tian, Jinguo Wang, Nitesh Kumar, Tianheng Han, Yoji Kobayashi, Ying Liu, Thomas E. Mallouk, and Moses H. W. Chan, “Observation of Superconductivity in Granular Bi Nanowires Fabricated by Electrodeposition,” Nano Letters 6, 2773 (2006).
- [45] N.N. Zhuravlev, A.A. Stepanova, N.I. Zyuzin, “On the Superconductivity of the Compound BiPt,” J. Exptl. Theoret. Phys. (U.S.S.R.) 37, 880-881 (1959).
- [46] Jing Wang, Xuliang Chen , Yonghui Zhou , Chao An, Ying Zhou, Chuanchuan Gu , Mingliang Tian, and Zhaorong Yang, “Pressure-induced superconductivity in trigonal layered PtBi2 with triply degenerate point fermions,” Phys. Rev. B 103, 014507 (2021).
- [47] C. Q. Xu, B. Li, M. R. van Delft, W. H. Jiao, W. Zhou, B. Qian, Nikolai D. Zhigadlo, Dong Qian, R. Sankar, N. E. Hussey, and Xiaofeng Xu, “Extreme magnetoresistance and pressure-induced superconductivity in the topological semimetal candidate YBi,” Phys. Rev. B 99, 024110 (2019).
- [48] H. Gu, F. Tang, Y.-R. Ruan, J.-M. Zhang, R.-J. Tang, W. Zhao, R. Zhao, L. Zhang, Z.-D. Han, B. Qian, X.-F. Jiang, and Y. Fang, “Pressure effect on the topologically nontrivial electronic state and transport of lutecium monobismuthide,” Phys. Rev. Materials 4, 124204 (2020).
- [49] E. R. Schemm, W. J. Gannon, C. M. Wishne, W. P. Halperin, and A. Kapitulnik, “Observation of broken time-reversal symmetry in the heavy-fermion superconductor UPt3,” Science 345, 190-193 (2014).
- [50] I. M. Hayes, D. S. Wei, T. Metz, J. Zhang, Y. S. Eo, S. Ran, S. R. Saha, J. Collini, N. P. Butch, D. F. Agterberg, A. Kapitulnik, and J. Paglione, “Multicomponent superconducting order parameter in UTe2,” Science 373, 797-801 (2021).
- [51] F. Hütt, A. Yaresko, M. B. Schilling, C. Shekhar, C. Felser, M. Dressel, and A. V. Pronin, “Linear-in-Frequency Optical Conductivity in GdPtBi due to Transitions near the Triple Points,” Phys. Rev. Lett. 121, 176601 (2018).
- [52] Troels A. Bojesen, Egor Babaev, and Asle Sudbø, ‘Time reversal symmetry breakdown in normal and superconducting states in frustrated three-band systems,” Phys. Rev. B 88, 220511(R) (2013).
- [53] Troels A. Bojesen, Egor Babaev, and Asle Sudbø, ‘Phase transitions and anomalous normal state in superconductors with broken time-reversal symmetry,” Phys. Rev. B 89, 104509 (2014).
- [54] Rafael M. Fernandes, Peter P. Orth, and Jörg Schmalian, “Intertwined Vestigial Order in Quantum Materials: Nematicity and Beyond,” Annual Review of Condensed Matter Physics 10, 133 (2019).
- [55] P. T. How and S. K. Yip, “Absence of Ginzburg-Landau mechanism for vestigial order in the normal phase above a two-component superconductor,” Phys. Rev. B 107, 104514 (2023).
- [56] C. Paulsen, D. J. Hykel, K. Hasselbach, and D. Aoki, “Observation of the Meissner-Ochsenfeld Effect and the Absence of the Meissner State in UCoGe,” Phys. Rev. Lett. 109, 237001 (2012).
- [57] Manuel Laubach, Johannes Reuther, Ronny Thomale, and Stephan Rachel, “Rashba spin-orbit coupling in the Kane-Mele-Hubbard model,” Phys. Rev. B 90, 165136 (2014).
- [58] Jason K. Kawasaki, Abhishek Sharan, Linda I. M. Johansson, Martin Hjort, Rainer Timm, Balasubramanian Thiagarajan, Brian D. Schultz, Anders Mikkelsen, Anderson Janotti, and Chris J. Palmstrøm, “A simple electron counting model for half-Heusler surfaces,” Sci. Adv.4,eaar5832(2018).
- [59] Jiwoong Kim, Kajetan M. Fijalkowski, Johannes Kleinlein, Claus Schumacher, Anastasios Markou, Charles Gould, Steffen Schreyeck, Claudia Felser, and Laurens W. Molenkamp, “Molecular beam epitaxy of a half-Heusler topological superconductor candidate YPtBi,” Phys. Rev. Materials 7, 024802 (2023).
- [60] G. Kresse and J. Hafner, “Ab initio molecular dynamics for liquid metals,” Phys. Rev. B 47, 558(R) (1993).